OpenStudy (migitmack):

Part 1: Solve each of the quadratic equations below. Show your work. x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

x^2 - 25 => x^2 - 5^2 therefore it = (x-5)(x+5) so, \[x=\pm5\] x^2-2x-15=0 => (x-5)(x+3), so, x = 5 and x=-3. each solution to each graph represents the x- intercepts alike: both have an x-intercept of +5, both are quadratic. different: the ratio of y values to x values are different excluding x=5

OpenStudy (migitmack):

did that say x=+5?, it looks a bit weird